Find the derivative of the following function. m(t) = - 41( 315 - 1) 8 m (t) =
Answer #1

triven that mit) = -4+ (3+ - 18 ti worot both sides differentiate mlt) d It -4+ (3ť f [ {fg} = fig + gf] - *{3221 +632-4 d